OpenStudy (anonymous):

Suppose every edge of a cube measures fifteen centimeters. What is the volume of the cube? two to the fifteenth power square centimeters three to the fifteenth power cubic centimeters fifteen to the second power square centimeters fifteen to the third power cubic centimeters

OpenStudy (lgbasallote):

volume = edge x edge x edge...let the calculator or the paper do the rest :DDD

OpenStudy (anonymous):

V= l * w * h Plug in and solve: V= 15 * 15 * 15 V= 15^3